A proton synchrotron accelerates protons to a kinetic energy of 533 GeV. At this energy, calculate (a) the Lorentz factor, (b) the speed parameter, and (c) the magnetic field for which the proton orbit has a radius of curvature of 764 m.
Answer #1

we know that kinetic energy of relativistic particle is

K.E = mc2 ( - 1)

or

= K.E / mc2 + 1

where mc2 is rest energy of proton = 938.257 MeV

= 533e9 / 938.257e6 + 1

= 569.1

----------------------------------------------

speed parameter is found as

= sqrt ( 1 - 1 / 2)

= sqrt ( 1 - 1 / 569.12)

= 0.99999845

-----------------------------------------------

we know for charged particle in field

qvB = mv2 / r

r = mv / qB

in relativistic terms

r = mv / qB

where v = *

so,

B = mc2 * sqrt (2 - 1) / qcr

B = 938.257e6 * 1.6e-19 * sqrt (569.12 - 1) / 3e8 * 1.6e-19 * 764

B = 2.329 T